OpenStudy (anonymous):

how do i use the half angle formula for a 22.5 degree angle for both cosine and sine

OpenStudy (anonymous):

hey liz..hi see cos^2 t = (1 + cos(2t))/2 if t = 22.5 cos^2 (22.5) = (1 - cos 45)/2

OpenStudy (anonymous):

22.5 = 45/2

OpenStudy (anonymous):

The formula is \[\cos^{2} \theta = (1 + \cos (2\theta))/2\]